If the correlation coefficient is 0.930, what is the unexplained variation?
A) 7% B) 93% C) 13.5% D) 86.5%

Answer by stanbon(75887)

If the correlation coefficient is 0.930, what is the unexplained variation?
A) 7% B) 93% C) 13.5% D) 86.5%
-----
R^2 = 0.93^2 = 0.865 ; this is the explained variation).
---
1-R^2 = 0.135 is the unexplained variation.
===================
Cheers,
Stan H.